# An efficient algorithm for $\mathcal{F}$-subgraph-free Edge Deletion on graphs having a product structure

### v1

@@ 1 Introduction
Several graph modification problems, including previous examples, are NP-complete [18, 19, 37, 53, 67]. Even when $\mathcal{F}$ consists of a single graph $F$, the NP-hardness of $\mathcal{F}$-subgraph-free Edge Deletion has been shown for several graphs $F$. For example, Anaso and Hirata [5] proved the NP-completeness when $F$ is 33-connected; Alon, Shapira, and Sudakov proved it when $F$ is not bipartite [2]; and Gishboliner, Levanzov, and Shapira proved it when $F$ is a forest that is not a star-forest [44]. Thus, exploring their complexity with additional constraints has emerged as an important research direction.
One direction is to design fixed-parameter tractable (FPT) algorithms, which are algorithms that run in time $f(\textsf{p})\cdot n^{O(1)}$ for some predetermined parameter p. A natural question in this direction is whether there exists an FPT algorithm when parameterized by the solution size $k$. For example, Pilipczuk, Pilipczuk, and Wrochna [62] proved that Edge Bipartization can be solved in time $O(1.977^{k})\cdot nm$ on $n$-vertex $m$-edge graphs. Moreover, if $\mathcal{F}$ is a fixed finite family of graphs, then there is a simple FPT algorithm for $\mathcal{F}$-subgraph-free Edge Deletion when parameterized by $k$, by using the bounded search tree technique [23]. However, even when $\mathcal{F}=\{F\}$, we cannot design an FPT algorithm for the same problem parameterized by $k+\lvert V(F)\rvert$ since this problem includes well-known $\textsf{W}[1]$-complete problems such as Clique and Independent Set [30], which are unlikely to have FPT algorithms unless the Exponential-Time Hypothesis (ETH) fails [23]. [^1] The Exponential-Time Hypothesis states that 33-SAT cannot be solved in time $2^{o(n)}$. In particular, ETH implies $\textsf{FPT}\neq\textsf{W}[1]$.
@@
Disk graphs have an involved structure in general, so it is natural to consider a parameter that measures their complexity. In this paper, we focus on the local radius, which is a parameter measuring the complexity of the disk graph and was recently introduced by Lokshtanov, Panolan, Saurabh, Xue, and Zehavi [58]. Roughly speaking, it measures the number of arcs we need to cross to connect two points in a disk. See Section 2 for the formal definition. A class of disk graphs with bounded local radius includes many interesting subclasses. For example, planar graphs are exactly the disk graphs of local radius 11 [57] and disk graphs of maximum degree at most $\Delta$ have local radius at most $2\Delta$. Lokshtanov, Panolan, Saurabh, Xue, and Zehavi [58] proved that a disk graph with a small local radius admits structural properties analogous to the planar graph, such as bidimensionality theory [26] or the Grid Minor Theorem [64]. Motivated by these results, we design an algorithm for $\mathcal{F}$-subgraph-free Edge Deletion on disk graphs of bounded local radius.
@@ 1.1 Product structure of graphs
Every planar graph is a subgraph of $H\boxtimes P$ for some graph $H$ of tree-width at most 66 and a path $P$.
@@
Given a planar graph $G$, there is a linear-time algorithm that finds a strong product $H\boxtimes P$ as well as a corresponding embedding of $G$ into $H\boxtimes P$, such that $H\boxtimes P$ contains $G$ as a subgraph, $P$ is a path, and the tree-width of $H$ is at most 66.
@@ 1.3 Related works
A contraction sequence of a trigraph $G$ is a sequence $G_{n},\ldots,G_{1}$ of trigraphs such that $G_{n}=G$, $G_{1}=K_{1}$, and $G_{i}$ is obtained from $G_{i+1}$ by contracting a pair of distinct vertices in $G_{i+1}$ for each ii. For an integer $d\geq 0$, a contraction sequence $G_{n},\ldots,G_{1}$ is a $d$-sequence if the graph $(V(G_{i}),R(G_{i}))$ has maximum degree at most $d$ for each $i\in[n]$. The twin-width of $G$, denoted by $\textsf{tww}(G)$, is the minimum integer $d$ such that a $d$-sequence of $G$ exists.
@@
Disk graphs exhibit a rich structure, leading to both positive and negative conclusions about the existence of a product structure. A positive answer comes from unit disk graphs, which are the intersection graphs of disks of radius 11 in the Euclidean plane. While the class of all unit disk graphs does not have a product structure due to the existence of large cliques, these are the only obstacles that prohibit the product structure. For each integer $t\geq 1$, the class of $K_{t}$-free unit disk graphs has a product structure [35].
@@ 2 Preliminaries
We refer to [27] for undefined notions. All graphs in this paper are finite and simple, that is, graphs have neither loops nor multiple edges. For a graph $G$, we denote by $V(G)$ and $E(G)$ the vertex set and the edge set of $G$, respectively. A graph $H$ is a subgraph of a graph $G$ if $V(H)\subseteq V(G)$ and $E(H)\subseteq E(G)$. A graph $G$ contains a graph $H$ if $G$ has a subgraph isomorphic to $H$; such a subgraph of $G$ is called a copy of $H$. Given a graph $H$, we say a graph $G$ is $H$-subgraph-free if $G$ has no copy of $H$. Given a family $\mathcal{F}$ of graphs, we also say that a graph $G$ is $\mathcal{F}$-subgraph-free if $G$ is $H$-subgraph-free for every $H\in\mathcal{F}$. For a set AA of vertices in a graph $G$, we denote by $G\setminus A$ the graph obtained from $G$ by removing all vertices in AA. We denote by $G[A]$ the induced subgraph of $G$ induced by AA, that is, $G[A]=G\setminus(V(G)\setminus A)$. For $v\in V(G)$, we also use the notation $G\setminus v$ to indicate $G\setminus\{v\}$. Similarly, for a set $S$ of edges in a graph $G$, we denote by GSG-S the graph obtained from $G$ by removing all edges in $S$.
@@ 4.1 Existence of the product structure
Due to Theorem 1.6, the arrangement graph $A_{\mathcal{D}}$ of $\mathcal{D}$ is a subgraph of $J\boxtimes P$ for some graph $J$ of tree-width at most 66 and a path $P$. Moreover, by Lemma 4.2, each vertex in $A_{\mathcal{D}}$ is contained in at most $2\rho+1$ distinct disks in $\mathcal{D}$. Let $\widetilde{A_{\mathcal{D}}}$ the graph obtained from $A_{\mathcal{D}}$ by applying a $(v,2\rho+1)$-blow-up for each $v\in V(A_{\mathcal{D}})$. Then $A_{\mathcal{D}}^{\prime}$ is a subgraph of $\widetilde{A_{\mathcal{D}}}$ and therefore, $A_{\mathcal{D}}^{\prime}\subseteq J\boxtimes P\boxtimes K_{2\rho+1}$ by 4.3. Thus, by Lemma 4.6 and Theorem 4.5, $G$ is a subgraph of $J^{\prime}\boxtimes P\boxtimes K_{(2\rho+1)^{3}}$ for some graph $J^{\prime}$ of tree-width at most ${2\rho+7\choose 6}-1$. Let $H\coloneqq J^{\prime}\boxtimes K_{(2\rho+1)^{3}}$. Then $G$ is a subgraph of $H\boxtimes P$, and $\mathsf{tw}(H)\leq\left({2\rho+7\choose 6}-1\right)\cdot(2\rho+1)^{3}=O(\rho^{9})$ by Lemma 3.1. ∎
@@ 5 Hardness for $\mathcal{F}$-subgraph-free edge deletion
For a clause $C_{i}\in\mathcal{C}$, let $x_{1}^{i},x_{2}^{i},x_{3}^{i}\in\mathcal{X}$ be its neighbors in $G_{\varphi}$, whose indices are assigned according to the cyclic order in the planar drawing of $G_{\varphi}$. We first define an auxiliary graph $W$ on 66 vertices as follows: We start with a cycle of length 66 on the vertex set $\mathinner{\left\{a_{1},a_{2},\ldots,a_{6}\right\}}$ and add three chords $a_{1}a_{5}$, $a_{2}a_{4}$, and $a_{2}a_{5}$.
@@
We first note that a triangle tiling of $W$ that covers $a_{4},a_{5},a_{6}$ must be a perfect triangle tiling. It is because $a_{4},a_{5},a_{6}$ do not form a triangle, so a triangle tiling that covers $a_{4},a_{5},a_{6}$ should contain at least 66 vertices. Also note that there is a unique such tiling that consists of $a_{1}a_{5}a_{6}$ and $a_{2}a_{3}a_{4}$.

### v2

@@ 1 Introduction
Several graph modification problems, including previous examples, are NP-complete [18, 19, 37, 53, 67]. Even when $\mathcal{F}$ consists of a single graph $F$, the NP-hardness of $\mathcal{F}$-subgraph-free Edge Deletion has been shown for several graphs $F$. For example, Anaso and Hirata [5] proved the NP-completeness when $F$ is $3$-connected; Alon, Shapira, and Sudakov proved it when $F$ is not bipartite [2]; and Gishboliner, Levanzov, and Shapira proved it when $F$ is a forest that is not a star-forest [44]. Thus, exploring their complexity with additional constraints has emerged as an important research direction.
One direction is to design fixed-parameter tractable (FPT) algorithms, which are algorithms that run in time $f(\textsf{p})\cdot n^{O(1)}$ for some predetermined parameter p. A natural question in this direction is whether there exists an FPT algorithm when parameterized by the solution size $k$. For example, Pilipczuk, Pilipczuk, and Wrochna [62] proved that Edge Bipartization can be solved in time $O(1.977^{k})\cdot nm$ on $n$-vertex $m$-edge graphs. Moreover, if $\mathcal{F}$ is a fixed finite family of graphs, then there is a simple FPT algorithm for $\mathcal{F}$-subgraph-free Edge Deletion when parameterized by $k$, by using the bounded search tree technique [23]. However, even when $\mathcal{F}=\{F\}$, we cannot design an FPT algorithm for the same problem parameterized by $k+\lvert V(F)\rvert$ since this problem includes well-known $\textsf{W}[1]$-complete problems such as Clique and Independent Set [30], which are unlikely to have FPT algorithms unless the Exponential-Time Hypothesis (ETH) fails [23]. [^1] The Exponential-Time Hypothesis states that $3$-SAT cannot be solved in time $2^{o(n)}$. In particular, ETH implies $\textsf{FPT}\neq\textsf{W}[1]$.
@@
Disk graphs have an involved structure in general, so it is natural to consider a parameter that measures their complexity. In this paper, we focus on the local radius, which is a parameter measuring the complexity of the disk graph and was recently introduced by Lokshtanov, Panolan, Saurabh, Xue, and Zehavi [58]. Roughly speaking, it measures the number of arcs we need to cross to connect two points in a disk. See Section 2 for the formal definition. A class of disk graphs with bounded local radius includes many interesting subclasses. For example, planar graphs are exactly the disk graphs of local radius $1$ [57] and disk graphs of maximum degree at most $\Delta$ have local radius at most $2\Delta$. Lokshtanov, Panolan, Saurabh, Xue, and Zehavi [58] proved that a disk graph with a small local radius admits structural properties analogous to the planar graph, such as bidimensionality theory [26] or the Grid Minor Theorem [64]. Motivated by these results, we design an algorithm for $\mathcal{F}$-subgraph-free Edge Deletion on disk graphs of bounded local radius.
@@ 1.1 Product structure of graphs
Every planar graph is a subgraph of $H\boxtimes P$ for some graph $H$ of tree-width at most $6$ and a path $P$.
@@
Given a planar graph $G$, there is a linear-time algorithm that finds a strong product $H\boxtimes P$ as well as a corresponding embedding of $G$ into $H\boxtimes P$, such that $H\boxtimes P$ contains $G$ as a subgraph, $P$ is a path, and the tree-width of $H$ is at most $6$.
@@ 1.3 Related works
A contraction sequence of a trigraph $G$ is a sequence $G_{n},\ldots,G_{1}$ of trigraphs such that $G_{n}=G$, $G_{1}=K_{1}$, and $G_{i}$ is obtained from $G_{i+1}$ by contracting a pair of distinct vertices in $G_{i+1}$ for each $i$. For an integer $d\geq 0$, a contraction sequence $G_{n},\ldots,G_{1}$ is a $d$-sequence if the graph $(V(G_{i}),R(G_{i}))$ has maximum degree at most $d$ for each $i\in[n]$. The twin-width of $G$, denoted by $\textsf{tww}(G)$, is the minimum integer $d$ such that a $d$-sequence of $G$ exists.
@@
Disk graphs exhibit a rich structure, leading to both positive and negative conclusions about the existence of a product structure. A positive answer comes from unit disk graphs, which are the intersection graphs of disks of radius $1$ in the Euclidean plane. While the class of all unit disk graphs does not have a product structure due to the existence of large cliques, these are the only obstacles that prohibit the product structure. For each integer $t\geq 1$, the class of $K_{t}$-free unit disk graphs has a product structure [35].
@@ 2 Preliminaries
We refer to [27] for undefined notions. All graphs in this paper are finite and simple, that is, graphs have neither loops nor multiple edges. For a graph $G$, we denote by $V(G)$ and $E(G)$ the vertex set and the edge set of $G$, respectively. A graph $H$ is a subgraph of a graph $G$ if $V(H)\subseteq V(G)$ and $E(H)\subseteq E(G)$. A graph $G$ contains a graph $H$ if $G$ has a subgraph isomorphic to $H$; such a subgraph of $G$ is called a copy of $H$. Given a graph $H$, we say a graph $G$ is $H$-subgraph-free if $G$ has no copy of $H$. Given a family $\mathcal{F}$ of graphs, we also say that a graph $G$ is $\mathcal{F}$-subgraph-free if $G$ is $H$-subgraph-free for every $H\in\mathcal{F}$. For a set $A$ of vertices in a graph $G$, we denote by $G\setminus A$ the graph obtained from $G$ by removing all vertices in $A$. We denote by $G[A]$ the induced subgraph of $G$ induced by $A$, that is, $G[A]=G\setminus(V(G)\setminus A)$. For $v\in V(G)$, we also use the notation $G\setminus v$ to indicate $G\setminus\{v\}$. Similarly, for a set $S$ of edges in a graph $G$, we denote by GSG-S the graph obtained from $G$ by removing all edges in $S$.
@@ 4.1 Existence of the product structure
Due to Theorem 1.6, the arrangement graph $A_{\mathcal{D}}$ of $\mathcal{D}$ is a subgraph of $J\boxtimes P$ for some graph $J$ of tree-width at most $6$ and a path $P$. Moreover, by Lemma 4.2, each vertex in $A_{\mathcal{D}}$ is contained in at most $2\rho+1$ distinct disks in $\mathcal{D}$. Let $\widetilde{A_{\mathcal{D}}}$ the graph obtained from $A_{\mathcal{D}}$ by applying a $(v,2\rho+1)$-blow-up for each $v\in V(A_{\mathcal{D}})$. Then $A_{\mathcal{D}}^{\prime}$ is a subgraph of $\widetilde{A_{\mathcal{D}}}$ and therefore, $A_{\mathcal{D}}^{\prime}\subseteq J\boxtimes P\boxtimes K_{2\rho+1}$ by 4.3. Thus, by Lemma 4.6 and Theorem 4.5, $G$ is a subgraph of $J^{\prime}\boxtimes P\boxtimes K_{(2\rho+1)^{3}}$ for some graph $J^{\prime}$ of tree-width at most ${2\rho+7\choose 6}-1$. Let $H\coloneqq J^{\prime}\boxtimes K_{(2\rho+1)^{3}}$. Then $G$ is a subgraph of $H\boxtimes P$, and $\mathsf{tw}(H)\leq\left({2\rho+7\choose 6}-1\right)\cdot(2\rho+1)^{3}=O(\rho^{9})$ by Lemma 3.1. ∎
@@ 5 Hardness for $\mathcal{F}$-subgraph-free edge deletion
For a clause $C_{i}\in\mathcal{C}$, let $x_{1}^{i},x_{2}^{i},x_{3}^{i}\in\mathcal{X}$ be its neighbors in $G_{\varphi}$, whose indices are assigned according to the cyclic order in the planar drawing of $G_{\varphi}$. We first define an auxiliary graph $W$ on $6$ vertices as follows: We start with a cycle of length $6$ on the vertex set $\mathinner{\left\{a_{1},a_{2},\ldots,a_{6}\right\}}$ and add three chords $a_{1}a_{5}$, $a_{2}a_{4}$, and $a_{2}a_{5}$.
@@
We first note that a triangle tiling of $W$ that covers $a_{4},a_{5},a_{6}$ must be a perfect triangle tiling. It is because $a_{4},a_{5},a_{6}$ do not form a triangle, so a triangle tiling that covers $a_{4},a_{5},a_{6}$ should contain at least $6$ vertices. Also note that there is a unique such tiling that consists of $a_{1}a_{5}a_{6}$ and $a_{2}a_{3}a_{4}$.